OpenStudy (anonymous):

Equation of a line that passes through the given point and is perpendicular to the given line in slope intercept form: y=x-3 (3,7)

OpenStudy (anonymous):

Use the point slope formula: y - y1 = m(x-x1). you have your point (x1, y1) = (3,7) all you need is your slope. Since the line you are looking for is perpendicular to the line given, its slope is the negative reciprocal of the slope of the line given. Use the calculated slope and the point in the point slope formula to solve for the equation of the line you are looking for

OpenStudy (anonymous):

so then the negative reciprocal of x will it be -1/1 ?? I'm not sure but I think it is

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

so the slope of the line you are looking for will be (-1/1) = -1

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh my bad it wouldnt end up as a fraction right?

OpenStudy (anonymous):

ok, so you are looking for a line that is perpendicular to the line y= x - 3. This line has a slope of +1. since you are looking for a line perpendicular to this line, its slope will be the negative reciprocal of + 1, which is -1.

OpenStudy (anonymous):

(-1/1) reduces to -1

OpenStudy (anonymous):

it could be left as a fraction because (-1/1) = -1, they are equivalent that's like using (5/1) instead of 5. They're virtually the same

OpenStudy (anonymous):

ok then thanks. so then after since its already in y incercept form i need to plug in the points given.after that it will give me b=10

OpenStudy (anonymous):

i think i got the final answer,is it y=-1x+10 ?

OpenStudy (anonymous):

yes, that's correct

OpenStudy (anonymous):

WHOOHOO! haha:p will you mind helping me in another problem?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

I'll put another post though!